2. If 12% of a population of parts is defective, what is the probability of randomly selecting 60 parts and finding that 8 or more are defective?


1
Expert's answer
2022-02-01T15:30:42-0500

"P=C_{60}^8\cdot 0.12^8\cdot 0.88^{52}=0.14=14\%."